# IC/DC: Surpassing Heuristic Solvers in Combinatorial Optimization with Diffusion Models

### v1

@@ A.3 Constraint Loss
$\mathcal{L}_{\text{cst}}(\theta)$ $=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{t}\sim q(X_{t}|X_{0})}\left[C(p_{\theta}(X_{0}|X_{t},c))\right],$
@@ A.4 Diffusion Loss
By combining reward-weighted diffusion loss A.2 and constraint loss A.3, our diffusion objective:
@@ B.1 Choice of Transition Matrices
According to (Austin et al. 2021), it is argued that incorporating domain-specific structures into the transition matrices $Q_{t}$ within the diffusion process is a reasonable approach. In our case, due to the inherent sparsity in the solution matrices of CO problems, the marginal distribution of feasible solutions significantly deviates from the uniform distribution commonly used in standard diffusion processes. Therefore, we design this noise transition matrix to align with the prior distribution of feasible solutions $\bar{q}$.
@@
In ATSP, when considering $|\mathcal{N}|$ cities, the solution involves travelling through all $|\mathcal{N}|$ cities. The prior distribution follows the following:
$\bar{q}=[1-\frac{1}{|\mathcal{N}|},\frac{1}{|\mathcal{N}|}]^{\top}$
@@
$\bar{q}=[1-\frac{1}{|\mathcal{J}|},\frac{1}{|\mathcal{J}|}]^{\top}$
@@ B.2 Problem Encoder
$S_{\text{inter}}$ $=\text{softmax}(AW_{\text{inter}}A^{\top})$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{A}|},$
$S_{\text{intra}}$ $=\text{ReLU}(AW_{\text{intra}}{B}^{\top})$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{B}|},$
$S$ $=\text{concat}(S_{\text{inter}}S_{\text{intra}},D)$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{B}|\times 2},$
$\tilde{D}$ $=\text{MLP}(S)$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{B}|},$
$\tilde{A}$ $=\text{softmax}(\tilde{D})BW_{\text{v}}$ $\in\mathbb{R}^{|\mathcal{A}|\times d},$
@@
$\hat{A}=\text{BN}(A+\tilde{A}),\quad A^{\prime}=\text{BN}(\hat{A}+\text{MLP}(\hat{A})),$
where BN refers to a batch normalization (Ioffe and Szegedy 2015). The process for updating $B$ to $B^{\prime}$ is computed in the same way.
@@ B.3 Denoiser
As illustrated in the bottom-left side of Figure 2, the denoiser consists of $L^{\prime}$ layers, where each layer gets input of $(A,B,X,t)$ and outputs $(A^{\prime},B^{\prime},X^{\prime})$, which is processed as follows:
$\hat{x}_{i,j}^{\ell+1}$ $=P^{\ell}x_{i,j}^{\ell}+Q^{\ell}h_{i}^{\ell}+R^{\ell}h_{j}^{\ell},$
$x_{i,j}^{\ell+1}$ $=x_{i,j}^{\ell}+\text{MLP}_{x}(\text{BN}(\hat{x}_{i,j}^{\ell+1}))+\text{MLP}_{t}(\mathbb{t}),$
$h_{i}^{\ell+1}$ $=h_{i}^{\ell}+\text{ReLU}\left(\text{BN}(U^{\ell}_{a}h_{i}^{\ell}+\sum_{j\in\mathcal{N}_{i}}(\sigma(\hat{x}_{i,j}^{\ell+1})\odot V^{\ell}_{b}h_{j}^{\ell}))\right),$
$h_{j}^{\ell+1}$ $=h_{j}^{\ell}+\text{ReLU}\left(\text{BN}(U^{\ell}_{b}h_{j}^{\ell}+\sum_{i\in\mathcal{N}_{j}}(\sigma(\hat{x}_{j,i}^{\ell+1})\odot V^{\ell}_{a}h_{i}^{\ell}))\right),$
@@
The matrices $U^{\ell}_{a},U^{\ell}_{b},V^{\ell}_{a},V^{\ell}_{b},P^{\ell},Q^{\ell},R^{\ell}\in\mathbb{R}^{d\times d}$ are learnable parameters of the $\ell$-th layer. SUM pooling is denoted by $\sum$ (Xu et al. 2018), the sigmoid function is represented by $\sigma$, and the Hadamard product is denoted by $\odot$. $\mathcal{N}_{i}$ denotes the neighborhood of node $i$ among the $B$ items, while $\mathcal{N}_{j}$ denotes the neighborhood of node $j$ among the $A$ items. Additionally, the variable $\mathbb{t}$ denotes the sinusoidal features (Vaswani et al. 2017) corresponding to the denoising timestep $t$. After the final layer $\mathfrak{L}$, the output $x^{\mathfrak{L}}_{i,j}$ is passed through a linear layer to obtain clean matrix $X_{0}=\{x_{0,0},...,x_{I,J}\}$.
@@ C.2 Baseline
Constraint Programming with Satisfiability (CP-SAT) is a highly efficient solver developed as part of the OR-Tools suite, designed for solving integer programming problems. It is particularly effective for solving scheduling problems, including PMSP. For PMSP our MIP model is based on Avalos-Rosales, Alvarez, and Angel-Bello (2013)
@@
$C_{j}:\text{Completion time of job j}$
$x_{ijk}:\begin{cases}1&\text{if job $k$ is processed directly after job $j$ on machine $i$}\\ 0&\text{otherwise}\end{cases}.$
$p_{ik}:\text{Time required to process job $j$ on machine $i$}$
$M:\text{Set of machines}$
$J:\text{Set of jobs to schedule}$
@@
$V:\text{A large positive number}$
@@
We utilize the implementations provided by Kwon et al. (Kwon et al. 2021) and follow their setting. Both GA and PSO run on GPUs.
@@
MatNet, proposed by (Kwon et al. 2021), adapts the attention model (Kool, Van Hoof, and Welling 2018) to be applicable to bipartite graphs. It employs cross-attention mechanisms in place of self-attention to facilitate message passing between the two set of nodes, thereby encoding relationship information more effectively. The model is trained using reinforcement learning. Solution are generated in an auto-regressive manner, where each node is sequentially selected based on the current state of the solution.
@@
We trained the diffusion model to estimate solution matrices using a supervised learning approach, similar to Difusco (Sun and Yang 2023). This method adapts graph-based denoising diffusion models to more naturally formulate combinatorial optimization problems and generate high-quality solutions. By explicitly modeling the node and edge selection process through corresponding random variables, the model effectively captures the problem’s structure. The model is trained through supervised learning.
@@
Similar to DDPO, as proposed by  (Black et al. 2023), which demonstrates that framing the denoising process as a multi-step decision-making problem enables policy gradient algorithms to directly optimize diffusion models for downstream objectives, we trained our diffusion model within the reinforcement learning framework. We used the REINFORCE algorithm (Williams 1992) to map the denoising process to the MDP framework.
@@ C.3 Training Detail
Following the implementation of (Kwon et al. 2021), we utilize 25 chromosomes with a mutation rate and crossover ratio both set to 0.3. Among the 25 initial chromosomes, one is initialized with the solution from the SJF heuristic. The best-performing chromosome is retained across all iterations. We run 1000 iterations per instance.
@@
Following the implementation of  (Kwon et al. 2021), we utilize 25 particles with an inertial weight of 0.7. Both the cognitive and social constants are set to 1.5. Additionally, one particle is initialized with the solution from the SJF heuristic. We run 1000 iterations per instance.
@@
We use the same hyperparameters reported in  (Kwon et al. 2021), with the exception that the number of stages is set to 1 instead of 3.
@@
Using CP-SAT, we generate 128,000 training samples for supervised learning. Since the solution matrices for PMSP are not square matrices, we use IC/DC’s encoder in combination with Difusco’s decoder, employing 12 layers for both encoder and the decoder. The denoising timestep is set to 1000 during training but reduced to 20 for faster inference. All other hyperparameters, follow those reported in  (Sun and Yang 2023).
@@
We train our diffusion model with the objective function
$\nabla_{\theta}\mathcal{J}_{\text{DDRL}}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}(X_{t-1}|X_{t},c)r(X_{0},c)\right]$ proposed by (Black et al. 2023). We use the same encoder and decoder architecture as IC/DC, consisting of 5 encoder layers and 2 decoder layers. Apart from setting the denoising timestep to 20, all other hyperparameters are consistent with those used in IC/DC (ours).
@@ D.1 Tmat class
While random distance matrices can be generated by choosing random integers, such matrices lack meaningful correlations between distances and doesn’t reflect practical scenarios. Instead we are interested in problems which ATSP instances have the triangle inequality so called ”Tmat class” (Kwon et al. 2021; Cirasella et al. 2001). That is, for a distance matrix $D$ with elements $d_{ij}$ representing the distance between city $c_{i}$ and $c_{j}$, if $d(c_{i},c_{j})\geq d(c_{i},c_{k})+d(c_{k},c_{j})$ then we set $d(c_{i},c_{j})=d(c_{i},c_{j})+d(c_{k},c_{j})$, while diagonal elements are maintained as $d(c_{i},c_{i})=0$. We repeat this procedure until no more changes can be made.
@@ D.2 Baseline
Mixed integer programming (MIP) is an optimization technique used to solve problems where some of the variables are required to be integers while others can be continuous. Solution methods such as branch and bound, branch and cut etc. are used to solve these kind of problems. To we use CPLEX (IBM 2022; Bliek1ú, Bonami, and Lodi 2014), one of the popular commercial optimization software use by the OR community and solve our test instances through benders decomposition (Rahmaniani et al. 2017). The MIP model serves as the mathematical representation of the problem. For ATSP, our MIP model is based on the formulation presented by Miller, Tucker, and Zemlin (1960).
@@
$i,j:\text{City index}$
$N:\text{Number of Cities}$
$c_{ij}:\text{Distance from city }i\text{ to city }j$
$x_{ij}:\begin{cases}1&\text{if you move from city $i$ to city $j$}\\ 0&\text{otherwise}\end{cases}.$
$u_{i}:\text{arbitrary numbers representing the order of city $i$in the tour}$
@@
As the name suggests, Nearest Neighbor (NN), Nearest Insertion (NI), and Furthest Insertion (FI) are straightforward simple greedy-selection algorithms frequently used as baselines for TSP algorithms. We use the implementations provided by  (Kwon et al. 2021) which are implemented in C++.
@@ D.3 Training Detail
We utilize the checkpoints provided by. (Kwon et al. 2021) and evaluate them on the same problem instances.
@@
Using LKH-3, we generate 128,000 training samples for supervised learning. We use Difusco’s encoder and decoder, employing 12 layers for both encoder and the decoder. The denoising timestep is set to 1000 during training but reduced to 20 for faster inference. All other hyperparameters, follow those reported in  (Sun and Yang 2023).
@@
We train our diffusion model with the objective function
$\nabla_{\theta}\mathcal{J}_{\text{DDRL}}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}(X_{t-1}|X_{t},c)r(X_{0},c)\right]$ proposed by (Black et al. 2023). We use the same encoder and decoder architecture as IC/DC, consisting of 5 encoder layers and 2 decoder layers. Apart from setting the denoising timestep to 20, all other hyperparameters are consistent with those used in IC/DC.
@@ D.4 Complex Feature Data
On the other hand, IC/DC is capable of handling these diverse types of data and demonstrates strong generalization performance. By the problem encoder, the information $R$ for each city is processed through an embedding layer with dimension $d$ and then sent to $A$ and $B$. The relational information between cities, $S$ and $F$, is input as $D=S||F\in\mathbb{R}^{|\mathcal{N}|\times|\mathcal{N}|\times 2}$. As shown in Table 3, IC/DC outperforms other baselines and demonstrates strong potential for generalizability. Additionally, an example of the NP is illustrated in Figure 5.

### v4

@@ A.3 Constraint Loss
$\mathcal{L}_{\text{cst}}(\theta)$ $=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{t}\sim q(X_{t}|X_{0})}\left[C(p_{\theta}(X_{0}|X_{t},c))\right],$ (43)
@@ A.4 Diffusion Loss
By combining reward-weighted diffusion loss Sec. A.2 and constraint loss Sec. A.3, our diffusion objective:
@@ B.1 Choice of Transition Matrices
According to [2], it is argued that incorporating domain-specific structures into the transition matrices $Q_{t}$ within the diffusion process is a reasonable approach. In our case, due to the inherent sparsity in the solution matrices of CO problems, the marginal distribution of feasible solutions significantly deviates from the uniform distribution commonly used in standard diffusion processes. Therefore, we design this noise transition matrix to align with the prior distribution of feasible solutions $\bar{q}$.
@@
In ATSP, when considering $|\mathcal{N}|$ and $N_{\text{col}}$ cells, the solution involves travelling through all $|\mathcal{N}|$ cities. The prior distribution follows the following:
$\bar{q}=[1-\frac{1}{|\mathcal{N}|},\frac{1}{|\mathcal{N}|}]^{\top}$ (47)
@@
$\bar{q}=[1-\frac{1}{|\mathcal{J}|},\frac{1}{|\mathcal{J}|}]^{\top}$ (48)
@@ B.2 Problem Encoder
$S_{\text{inter}}$ $=\text{softmax}(AW_{\text{inter}}A^{\top})$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{A}|},$ (49)
$S_{\text{intra}}$ $=\text{ReLU}(AW_{\text{intra}}{B}^{\top})$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{B}|},$ (50)
$S$ $=\text{concat}(S_{\text{inter}}S_{\text{intra}},D)$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{B}|\times 2},$ (51)
$\tilde{D}$ $=\text{MLP}(S)$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{B}|},$ (52)
$\tilde{A}$ $=\text{softmax}(\tilde{D})BW_{\text{v}}$ $\in\mathbb{R}^{|\mathcal{A}|\times d},$ (53)
@@
$\hat{A}=\text{BN}(A+\tilde{A}),\quad A^{\prime}=\text{BN}(\hat{A}+\text{MLP}(\hat{A})),$ (54)
where BN refers to a batch normalization [13]. The process for updating $B$ to $B^{\prime}$ is computed in the same way.
@@ B.3 Denoiser
As illustrated in the bottom-left side of  Fig. 2, the denoiser consists of $L^{\prime}$ layers, where each layer gets input of $(A,B,X,t)$ and outputs $(A^{\prime},B^{\prime},X^{\prime})$, which is processed as follows:
$\hat{x}_{i,j}^{\ell+1}$ $=P^{\ell}x_{i,j}^{\ell}+Q^{\ell}h_{i}^{\ell}+R^{\ell}h_{j}^{\ell},$ (55)
$x_{i,j}^{\ell+1}$ $=x_{i,j}^{\ell}+\text{MLP}_{x}(\text{BN}(\hat{x}_{i,j}^{\ell+1}))+\text{MLP}_{t}(\mathbb{t}),$ (56)
$h_{i}^{\ell+1}$ $=h_{i}^{\ell}+\text{ReLU}\left(\text{BN}(U^{\ell}_{a}h_{i}^{\ell}+\sum_{j\in\mathcal{N}_{i}}(\sigma(\hat{x}_{i,j}^{\ell+1})\odot V^{\ell}_{b}h_{j}^{\ell}))\right),$ (57)
$h_{j}^{\ell+1}$ $=h_{j}^{\ell}+\text{ReLU}\left(\text{BN}(U^{\ell}_{b}h_{j}^{\ell}+\sum_{i\in\mathcal{N}_{j}}(\sigma(\hat{x}_{j,i}^{\ell+1})\odot V^{\ell}_{a}h_{i}^{\ell}))\right),$ (58)
@@
The matrices $U^{\ell}_{a},U^{\ell}_{b},V^{\ell}_{a},V^{\ell}_{b},P^{\ell},Q^{\ell},R^{\ell}\in\mathbb{R}^{d\times d}$ are learnable parameters of the $\ell$-th layer. SUM pooling is denoted by $\sum$ [31], the sigmoid function is represented by $\sigma$, and the Hadamard product is denoted by $\odot$. $\mathcal{N}_{i}$ denotes the neighborhood of node $i$ among the $B$ items, while $\mathcal{N}_{j}$ denotes the neighborhood of node $j$ among the $A$ items. Additionally, the variable $\mathbb{t}$ denotes the sinusoidal features [27] corresponding to the denoising timestep $t$. After the final layer $\mathfrak{L}$, the output $x^{\mathfrak{L}}_{i,j}$ is passed through a linear layer to obtain clean matrix $X_{0}=\{x_{0,0},...,x_{I,J}\}$.
@@ C.2 Baseline
Constraint Programming with Satisfiability (CP-SAT) is a highly efficient solver developed as part of the OR-Tools suite, designed for solving integer programming problems. It is particularly effective for solving scheduling problems, including PMSP. For PMSP our MIP model is based on Avalos-Rosales et al. [3]
@@
$C_{j}:\text{Completion time of job j}$ (69)
$x_{ijk}:\begin{cases}1&\text{if job $k$ is processed directly after job $j$ on machine $i$}\\ 0&\text{otherwise}\end{cases}.$ (70)
$p_{ik}:\text{Time required to process job $j$ on machine $i$}$ (71)
$M:\text{Set of machines}$ (72)
$J:\text{Set of jobs to schedule}$ (73)
@@
$V:\text{A large positive number}$ (75)
@@
We utilize the implementations provided by Kwon et al. [19] and follow their setting. Both GA and PSO run on GPUs.
@@
MatNet, proposed by [19], adapts the attention model [17] to be applicable to bipartite graphs. It employs cross-attention mechanisms in place of self-attention to facilitate message passing between the two set of nodes, thereby encoding relationship information more effectively. The model is trained using reinforcement learning. Solution are generated in an auto-regressive manner, where each node is sequentially selected based on the current state of the solution.
@@
We trained the diffusion model to estimate solution matrices using a supervised learning approach, similar to Difusco [26]. This method adapts graph-based denoising diffusion models to more naturally formulate combinatorial optimization problems and generate high-quality solutions. By explicitly modeling the node and edge selection process through corresponding random variables, the model effectively captures the problem’s structure. The model is trained through supervised learning.
@@
Similar to DDPO, as proposed by  [5], which demonstrates that framing the denoising process as a multi-step decision-making problem enables policy gradient algorithms to directly optimize diffusion models for downstream objectives, we trained our diffusion model within the reinforcement learning framework. We used the REINFORCE algorithm [30] to map the denoising process to the MDP framework.
@@ C.3 Training Detail
(Experimental detail) We evaluated the baselines and the proposed algorithm using two Intel Xeon Gold 6330 CPUs and an RTX 3090 GPU for both PMSP and ATSP.
@@
Following the implementation of [19], we utilize 25 chromosomes with a mutation rate and crossover ratio both set to 0.3. Among the 25 initial chromosomes, one is initialized with the solution from the SJF heuristic. The best-performing chromosome is retained across all iterations. We run 1000 iterations per instance.
@@
Following the implementation of  [19], we utilize 25 particles with an inertial weight of 0.7. Both the cognitive and social constants are set to 1.5. Additionally, one particle is initialized with the solution from the SJF heuristic. We run 1000 iterations per instance.
@@
We use the same hyperparameters reported in  [19], with the exception that the number of stages is set to 1 instead of 3.
@@
Using CP-SAT, we generate 128,000 training samples for supervised learning. Since the solution matrices for PMSP are not square matrices, we use IC/DC’s encoder, employing 5 layers. The denoising timestep is set to 20. To generate a feasible solution, we apply greedy decoding after the denoising process. The greedy decoding selects the maximum probability for each job based on a heatmap generated by the model. If there is more than one machine with the maximum probability, one of them is chosen randomly. All other hyperparameters, follow those reported in  [26]
@@
We train standard discrete diffusion model with the objective function
$\nabla_{\theta}\mathcal{J}_{\text{DDRL}}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}(X_{t-1}|X_{t},c)r(X_{0},c)\right]$ proposed by [5].We use the same encoder architecture as IC/DC, along with a GAT-based decoder, consisting of 5 encoder layers and 2 decoder layers.Additionally, we set the denoising timestep to 20. To enforce the model to output feasible solution, we apply greedy decoding, which is the same as the one used in Diffusion (SL). All other hyperparameters are consistent with those used in IC/DC.
@@ D.1 Tmat class
While random distance matrices can be generated by choosing random integers, such matrices lack meaningful correlations between distances and doesn’t reflect practical scenarios. Instead we are interested in problems which ATSP instances have the triangle inequality so called ”Tmat class” [19, 8]. That is, for a distance matrix $D$ with elements $d_{ij}$ representing the distance between city $c_{i}$ and $c_{j}$, if $d(c_{i},c_{j})\geq d(c_{i},c_{k})+d(c_{k},c_{j})$ then we set $d(c_{i},c_{j})=d(c_{i},c_{j})+d(c_{k},c_{j})$, while diagonal elements are maintained as $d(c_{i},c_{i})=0$. We repeat this procedure until no more changes can be made.
@@ D.2 Baseline
Mixed integer programming (MIP) is an optimization technique used to solve problems where some of the variables are required to be integers while others can be continuous. Solution methods such as branch and bound, branch and cut etc. are used to solve these kind of problems. To we use CPLEX [12, 6], one of the popular commercial optimization software use by the OR community and solve our test instances through benders decomposition [23]. The MIP model serves as the mathematical representation of the problem. For ATSP, our MIP model is based on the formulation presented by Miller et al. [20].
@@
$i,j:\text{City index}$ (80)
$N:\text{Number of Cities}$ (81)
$c_{ij}:\text{Distance from city }i\text{ to city }j$ (82)
$x_{ij}:\begin{cases}1&\text{if you move from city $i$ to city $j$}\\ 0&\text{otherwise}\end{cases}.$ (83)
$u_{i}:\text{arbitrary numbers representing the order of city $i$in the tour}$ (84)
@@
As the name suggests, Nearest Neighbor (NN), Nearest Insertion (NI), and Furthest Insertion (FI) are straightforward simple greedy-selection algorithms frequently used as baselines for TSP algorithms. We use the implementations provided by  [19] which are implemented in C++.
@@ D.3 Training Detail
We utilize the checkpoints provided by. [19] and evaluate them on the same problem instances.
@@
Using LKH-3, we generate 128,000 training samples for supervised learning. We use DIFUSCO’s encoder, employing 5 layers for encoder. We set the denoising timestep to 20. To enforce feasible solution, we apply greedy decoding to the heatmap generated after the denoising process. The greedy decoding starts at node 0 and chooses the maximum probability node among those not visited. All other hyperparameters, follow those reported in  [26].
@@
We train a standard discrete diffusion model with the objective function
$\nabla_{\theta}\mathcal{J}_{\text{DDRL}}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}(X_{t-1}|X_{t},c)r(X_{0},c)\right]$ proposed by [5]. We use the same encoder architecture as IC/DC, along with a GAT-based decoder, consisting of 5 encoder layers and 2 decoder layers. Additionally, we set the denoising timestep to 20. To enforce the model to output feasible solution, we apply greedy decoding, which is the same as the one used in Diffusion (SL). All other hyperparameters are consistent with those used in IC/DC.
@@ D.4 Complex Feature Data
On the other hand, IC/DC is capable of handling these diverse types of data and demonstrates strong generalization performance. By the problem encoder, the information $R$ for each city is processed through an embedding layer with dimension $d$ and then sent to $A$ and $B$. The relational information between cities, $S$ and $F$, is input as $D=S||F\in\mathbb{R}^{|\mathcal{N}|\times|\mathcal{N}|\times 2}$. As shown in Tab. 3, IC/DC outperforms other baselines and demonstrates strong potential for generalizability. Additionally, an example of the NP is illustrated in Fig. 6.
@@
Appendix E Additional experiment results
E.1 Distribution generalization full result
MatNet Diffusion (SL) Diffusion (RL) IC/DC
Node Tour Length $\downarrow$ Gap $\downarrow$ Tour Length $\downarrow$ Gap $\downarrow$ Tour Length $\downarrow$ Gap $\downarrow$ Tour Length $\downarrow$ Gap $\downarrow$
10 1.498 0.064% 1.744 15.213% 2.403 46.427% 1.497 0%
15 1.590 4.009% 1.765 14.411% 2.815 59.275% 1.528 0%
20* 1.534 0.013% 1.755 13.438% 3.331 73.875% 1.534 0%
25 2.006 25.712% 1.934 22.147% 3.328 72.986% 1.553 0.255%
30 2.243 36.678% 2.069 28.803% 3.501 77.375% 1.607 3.733%
35 2.462 45.888% 2.213 35.647% 3.678 81.762% 1.765 13.393%
40 2.640 51.849% 2.389 42.407% 3.800 83.971% 1.976 23.987%
45 2.798 57.116% 2.533 47.843% 3.936 86.743% 2.174 33.227%
50 2.950 62.150% 2.676 53.233% 4.090 90.010% 2.421 43.797%
Table 4: Full result on distribution generalization evaluation across different methods. * denotes the training distribution (20 nodes).
Tab. 4 shows the generalization performance of various models, including MatNet, Diffusion (SL), Diffusion (RL), and IC/DC, evaluated on node sizes from 10 to 50. Notably, IC/DC, trained on 20-node distributions, achieves a perfect gap of 0% at its training distribution and smaller node sizes (10 and 15 nodes), outperforming all other models.
As node size increase beyond the training distribution, IC/DC continues to exhibit strong generalization capabilities. At 30 nodes, IC/DC achieves a gap of 3.733%, outperforming MatNet (36.678%) and Diffusion models (SL: 28.803%, RL: 77.375%). Even at 50 nodes, IC/DC maintains its superiority with a gap of 43.797%, significantly better than MatNet (62.150%) and Diffusion models (SL: 53.233%, RL: 90.010%).
In summary, IC/DC consistently delivers superior results across various node sizes, demonstrating excellent adaptability and generalization. Its ability to maintain competitive performance, even on test distributions far from its training distribution, highlights its efficiency and robustness in CO problems.
Figure 7: Full results on optimality gap and inference time as a function of sampling step count in the parallel machine scheduling problem (PMSP). The left two plots correspond to a 3 machines & 50 jobs setup, while the right two plots represent a 5 machines & 50 jobs setup.
E.2 Reducing sampling steps for faster inference additional results
Fig. 7 demonstrates the relationship between varying sampling steps and their impact on the optimality gap and inference time. The results are shown for IC/DC and MatNet when solving PMSP instances with 20 jobs on 5 machines and 50 jobs on 3 machines. While neither IC/DC nor MatNet achieves a 0% optimality gap, IC/DC performs better in terms of optimality gap compared to MatNet at $T_{I}=T$.
The larger the number of samples during inference, the less the optimality gap is affected when reducing sampling steps ($T_{I}$). IC/DC outperforms MatNet at $T_{I}=\frac{1}{3}T$ and $T_{I}=\frac{1}{4}T$. Moreover, IC/DC achieves superior performance while maintaining competitive inference speed.